Circular No. 6508

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bv IN UGC 3432 Weidong Li, Yulei Qiu, Qiran Qiao, and Jingyao Hu, Beijing Astronom-

ical Observatory (BAO), on behalf of the BAO Supernova Survey, report their discovery of a supernova on unfiltered CCD images taken with the BAO 0.60-m reflector. SN 1996bv is located 3′′.6 northwest of the center of UGC 3432 (α = 6^{h}16^{m}.3, δ = +57^{o}03′, equinox 2000.0), and was estimated at the following magnitudes: Nov. 3 UT, 16.6; 12, 15.7; 16, 15.6. A V CCD image of the same field taken on Oct. 18 shows no star at this position. A low-resolution spectrum (0.96 nm/pixel, range 300–900 nm) taken with the BAO 2.16-m telescope by Li, Qiao, and J. Hao on Nov. 21.83 shows conspicuous Si II 635-nm absorption at 615 nm and typical features of type-Ia supernovae at maximum light (Si II, S II, Ca II, Mg II, etc.), with the overall spectrum resembling that of SN 1989B at maximum light.

GX 13+1 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, reports on behalf of the RXTE ASM team at GSFC and the Massachusetts Institute of Technology: “Analysis of the RXTE ASM lightcurve of the low-mass x-ray binary GX 13+1 (= 4U 1811– 17) obtained between 1996 Feb. 23 and Nov. 14 suggests the presence of a modulation with a period of 24.7 ± 1 days. The folded lightcurve on this period shows a mean peak-to-peak modulation of ∼ 15 percent. This periodicity is apparently confirmed by the Ariel 5 ASM lightcurve of GX 13+1, as a Fourier Transform of this shows that the largest peak, for periods < 125 days, occurs at 25.2 ± 0.15 days, consistent with the RXTE result. In the RXTE data, the modulation is most pronounced at low energies (1.3–4.8 keV) and is reduced at higher energies (4.8–12.2 keV). In addition, there is an anticorrelation between hardness ratio and intensity in the RXTE data, which may suggest a similarity with Cyg X-2 (Wijnands et al. 1996, Ap.J. 473, L45). If this periodicity is a persistent property of this source, it may also be present in the variability of the infrared counterpart.”

COMET 91P/RUSSELL 3 CCD magnitude estimates by J. V. Scotti, University of Arizona, with

the Spacewatch 0.9-m telescope: Nov. 18.53 UT, 21.0; 19.54, 21.4. The comet is of stellar appearance.

1996 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green